### v1

# Universal Omni-Language Models

Abstract
Recent advancements in multimodal large language models (MLLMs) have aimed to integrate and interpret data across diverse modalities. However, the capacity of these models to concurrently process and reason about multiple modalities remains inadequately explored, partly due to the lack of comprehensive modality-wise benchmarks. We introduce OmniBench, a novel benchmark designed to rigorously evaluate models’ ability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously. We define models capable of such tri-modal processing as omni-language models (OLMs). OmniBench is distinguished by high-quality human annotations, ensuring that accurate responses require integrated understanding and reasoning across all three modalities. Our main findings reveal that: i) open-source OLMs exhibit critical limitations in instruction-following and reasoning capabilities within tri-modal contexts; and ii) most baselines models perform poorly (below 50% accuracy) even when provided with alternative textual representations of images or/and audio. These results suggest that the ability to construct a consistent context from text, image, and audio is often overlooked in existing MLLM training paradigms. We advocate for future research to focus on developing more robust tri-modal integration techniques and training strategies to enhance OLM performance across diverse modalities. Codes and live leaderboard could be found at https://m-a-p.ai/OmniBench.
1 Introduction
The rapid advancement of artificial intelligence has ushered in a new era of multimodal large language models (MLLMs), capable of processing and interpreting diverse data types mainly including images, audio, and text (Li et al., 2024b). These models aim to emulate human-like understanding of the world by integrating information across multiple sensory modalities and learning a comprehensive context from the environment. While significant strides have been made in developing MLLMs that can handle two of the modalities, the ability to concurrently process and reason about the three aforementioned modalities remains a frontier yet to be fully explored.
The challenge in advancing MLLMs lies not only in their development but also in our capacity to evaluate their performance comprehensively. Current benchmarks often solely focus on image or audios, or limited image-text (Yue et al., 2024; Zhang et al., 2024) or audio-text combinations (Yang et al., 2024) for the dual-modality vision-language models (VLMs) (Laurençon et al., 2024) or audio-language models (ALMs) (Chu et al., 2023a; Deng et al., 2023). This gap in evaluation tools has hindered the community to assess and improve the holistic capabilities of models right before the dawn of general-purpose MLLMs.
To address this critical need, we introduce OmniBench, a pioneering universal multimodal benchmark designed to rigorously evaluate MLLMs’ capability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously, which we define as the omni-understanding and reasoning ability of the omni-language models (OLMs) (Sun et al., 2024; Zhan et al., 2024; Lu et al., 2024b). For instance, one can only derive the correct answer of the sample question in Figure 1 by: 1)))) recognizing elements from the given image and audio to reconstruct the context; 2)))) interpreting the semantics and relationships among the multimodal objects according to the textual instruction formed as question and options; 3)))) reasoning and then answering with the complementary information from all the modalities. We distinguishes OmniBench by enforcing a unique constraint: accurate responses necessitate an integrated understanding and reasoning of all multimodal contexts. This approach ensures a more realistic and challenging assessment of multimodal large language models, mirroring the complex, interconnected nature of human cognition. To ensure the evaluation reliability, the development of OmniBench relies on high-quality human annotations. Furthermore, OmniBench additionally includes the answer rationales provided by the annotators to enhance the validity and ensure the benchmark aligned with human-level understanding.
Our initial findings using OmniBench reveal critical limitations in the omni-understanding capabilities of existing MLLMs:
  • Although the existing open-source omni-language models have been trained with data in the three modalities, they could barely exceed the performance of random guess accuracy or even hard to follow the instruction when provided with image and audio together.
  • In the setting of using text as the alternatives of audios and images, the open-source VLMs and ALMs show a relatively better results but remain in a preliminary level of capability to understand the given tri-modality context.
  • Compared to the open-source OLMs, the proprietary model seems not suffering from performance drop when it comes to the full evaluation setting of our benchmark.
These results underscore the importance of OmniBench as a tool for identifying areas of improvement and guiding research in multimodal systems. In the following sections, we detail the data collection protocol of OmniBench, present our evaluation results on current state-of-the-art MLLMs, and discuss the implications of our findings for the future of research and development. Through OmniBench, we aim to catalyze advancements in MLLMs, pushing the boundaries of artificial intelligence towards true omni-understanding capabilities.
Figure 1: Example Data from Different Categories. The main contextual information is provided by the corresponding image and audio, while the question and options are expressed with text. Playable demos could be found at the homepage.
2 Related Work
Multimodal Large Language Models.
Many studies have been introduced to enable model to perceive complex audio signals and generate text responses based on human instructions. Notable examples in audio perception include SALMONN Tang et al. (2023), BLSPN Wang et al. (2023a), Speech-LLaMAN Wu et al. (2023), and Qwen-Audio Chu et al. (2023b), all of which demonstrate promising capabilities in audio-focused dialogues. Regarding the visual modality, the large visual language model also makes excellent achievements. BLIP2 Li et al. (2023) uses a Q-Former to align the visual knowledge with the textual information during the pre-training phase. LLaVA Liu et al. (2024b) is pre-trained on the GPT-4 generated multimodal language-image instruction-following data and uses a projection to fuse the vision module and the LM. LLaVA-Next Liu et al. (2024a) improves the single-image performance with a cost of an increasing number of image tokens. Based on the framework LLaVA mode, QwenVL Bai et al. (2023), CogVLM Wang et al. (2023b), and Yi-VL Young et al. (2024) all achieve significant success through a huge number of pre-training data. However, most existing multi-modal language models only focus on a certain modal and generate textual responses. While some models can process textual, audio, and visual inputs simultaneously, open-source models in this field exhibit relatively average capabilities compared to their closed-source counterparts. In the following context of this literature, we refer to the omni-language models (OLMs) as those capable of processing at least three different modalities of data simultaneously [^1] At the moment, we target the models concurrently processing with image, audio, and text as a starting point since these are the most well-explored modalities in the field..
Multimodal Understanding Benchmark.
The vision-language benchmarks aim to test models’ ability to combine visual and language data in tasks like OCR, spatial awareness, and multimodal reasoning skills. MM-Vet Yu et al. (2023) focuses on visual question answering (VQA), requiring models to interpret visual data and respond to queries. MMBench Liu et al. (2023b) evaluates models via multiple-choice tasks in both Chinese and English, covering diverse domains. MMStar (Chen et al., 2024a) conducts multi-task evaluations to test multimodal fusion capabilities. MMMU Yue et al. (2024) and CMMMU Zhang et al. (2024) assess model performance on complex vision-language tasks, emphasizing sophisticated multimodal reasoning. In addition, there are several audio-understanding benchmarks. Aishell1 Bu et al. (2017), Aishell2 Du et al. (2018), and Librispeech Panayotov et al. (2015) are designed for automatic speech recognition, while ClothoAQA targets audio QA tasks. For automatic audio captioning and vocal sound classification, researchers have curated Clotho Drossos et al. (2020) and VocalSound Gong et al. (2022). However, there is a significant lack of comprehensive understanding benchmarks to assess models’ ability to simultaneously process various modalities, including textual, audio, and visual inputs.
3 OmniBench
The OmniBench aims to create the first comprehensive benchmark for evaluating multimodal large language models that support simultaneous image, audio, and text inputs. While OmniBench is designed to evaluate the understanding capability of MLLMs on cross-modality complementary information, the models are required to interpret the multimodal input and provide accurate text answer. The problem could be formulated as following: given a tuple of (image, audio, text), the model is required to recognize the objects, re-build the contexts, and conduct reasoning based on the given information. The design logic and statistics of the dataset and the annotation protocols are introduced in this section.
Table 1: The Statistics of OmniBench Across Task Types. The word lengths of four options for each question are first averaged, and then the averages are calculated in group.
Statistic Action & Activity Story Description Plot Inference Object Identification & Description Contextual & Environmental Identity & Relationship Text & Symbols Count & Quantity Overall
Quantity 251 237 230 211 141 32 25 15 1142
Word Length
Question 4.68 5.75 7.47 7.00 6.85 6.22 7.32 8.72 6.25
Option 8.85 7.77 8.92 6.47 5.68 10.38 11.22 6.60 8.81
Img. Rationale 18.27 19.62 24.40 24.94 18.34 22.69 24.80 29.16 21.19
Audio Rationale 23.11 20.50 24.40 20.97 18.27 24.92 23.10 53.84 22.90
Audio Content 13.21 17.91 29.87 28.03 14.41 19.01 23.31 35.16 18.37
Multimodal Information
Img. Width 1283.75 1291.60 2394.93 1430.03 1141.39 1395.53 1338.51 1787.36 1322.36
Img. Height 842.32 776.11 2089.93 799.47 728.06 840.15 761.58 1168.04 818.64
Audio Len. (s) 7.35 9.82 11.22 11.43 8.03 8.63 11.43 15.63 9.22
Figure 2: The Data Distribution of OmniBench. Each chart represents the data samples grouped by task types and distinguished by audio types.
3.1 Benchmark Design
Building upon existing multimodal benchmarks, we propose a novel task type categorization in OmniBench that assesses a broad spectrum of reasoning and cognitive abilities. Our taxonomy progresses from fundamental perception (Object Identification & Description) to complex inference (Contextual & Environmental, Identity & Relationship). It incorporates temporal and logical order understanding of events (Action & Activity, Story Description, Plot Inference), spatial awareness (Contextual & Environmental), entity recognition (Object Identification & Description), symbolic processing (Text & Symbols), and quantitative reasoning (Count & Quantity). This comprehensive design evaluates both low-level perceptual skills and high-level cognitive functions, enabling a holistic assessment of multimodal language models’ (MLLMs) capabilities to recognize, describe, integrate information, understand context, and make nuanced inferences. OmniBench comprises 1142 question-answer pairs, with task type distribution, text length, and image and audio characteristics detailed in Table 1. The dataset’s audio content falls into three categories: speech (human vocal communication), sound events (non-speech natural, environmental and mechanical sounds), and music (various compositions and performances). Figure 2 illustrates the distribution of these audio types across task categories. This diverse task set aims to reveal the strengths and weaknesses of existing models in a tri-modal setting, potentially guiding the development of future omni-language models.
Figure 3: OmniBench Annotation Scheme. The example shows flawed data that didn’t pass inspection because the information in audio alone is sufficient to answer.
3.2 Annotation Protocol
Annotation Scheme.
Our annotation scheme is built upon a fundamental principle: the correct answer to each question must require information from both the image and audio components. This ensures that the benchmark effectively evaluates the model’s ability to analyze information across modalities. As shown in Figure 3, we implemented a rigorous annotation pipeline consisting of three stages: initial annotation, human inspection, and model inspection. Data samples that failed to meet our criteria at any stage were returned to annotators for revision, ensuring high-quality, multimodal-dependent samples. Through the whole process, $16$ annotators and $5$ quality inspectors are involved, all of which are full time industrial data annotation employee with higher education backgrounds.
The questions are formalized as multi-choice question-answering (MCQ) but try to maintain a consistent logic that suggests the only one possible and accurate answer, i.e., they can be potentially further re-organized into blank filling questions. Furthermore, when constructing the options, the annotators need to ensure at least one confusing wrong option. To ensure question difficulty, annotators were required to verify that questions and options were not trivially easy, lacked distinguishable patterns, and could not be answered by state-of-the-art MLLMs using image information alone. GPT-4 are allowed to use to provide initial annotator self-assessments of question quality. We restrict the images with a minimum resolution of 480P (854x480 pixels) and audio clips with a maximum duration of 30 seconds.
We implemented strict measures to maintain diversity across the dataset. This includes varying image and audio sources, limiting the frequency of individual speakers in audio clips to no more than five occurrences, and restricting the replication of similar instructions or questions. For instance, questions about specific environmental contexts were limited up to three samples. Importantly, annotators were required to provide rationales for correct answers, detailing the specific information that should be derived from the image and audio modalities respectively. This approach not only aided in quality inspection but also laid the groundwork for future fine-grained evaluation.
Figure 4: The Distribution of Inspection Frequency of the Passed Samples in OmniBench.
Quality Control.
Our quality control process was two-fold, including human inspection round and automatic inspection round assisted by MLLM. First, all annotated instruction-response pairs undergo cross-inspection by human annotators. Inspectors provide detailed reasons for any samples that failed to meet our stringent criteria, allowing for targeted revisions. Samples that pass human inspection are then subjected to a secondary check using a vision-language model LLaVA-1.6-34B (Liu et al., 2024a), where the the automatic quality inspection model is selected by considering trade-off between efficiency and performance. This automated process evaluates each sample under various ablation settings: image and text only, audio transcript and text only, and text only (repeated three times). Samples are only accepted if the model either rejected the task or made mistakes under these limited-information scenarios, confirming the necessity of both visual and auditory information for correct responses. We plot the distribution of the inspection frequency of the passed samples in Figure 4, where we could find that $76\%$ (868) of the passed samples do not require further modification under a well-defined annotation framework and $21.1\%$ of them requiring 1-2 times of revision. During the iterative quality checking, $9.58\%$ ($121$) QA pairs are defined as ”hard to recycle” and dumped after revisions and discussions.
4 Experiment Settings
Baseline Models.
We select three groups of MLLM baselines according to the modalities available: i. omni-language models: AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b); ii. vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series team (2024), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024a), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); iii. audio-language models: LTU series Gong et al. (2023), Mu-LLaMA (Liu et al., 2023a), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not categorize them into the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed [^2] The authors conclude from an investigation on September 22, 2024..
Omni-Understanding Evaluation.
The main focus of OmniBench is to evaluate how well could the MLLMs understand and reconstruct the context given information from image, audio and text modalities. Setting up questions with four available options for the models, we use accuracy, i.e., the ratio matched letter of the correct option and model response, as the evaluation metric (n.b., the accuracy of a random guess model is $25\%$ under this setting). Additionally, we test the models in an ablation setting of removing one of the image or audio inputs to further reveal a more comprehensive reasoning capability of the baselines and verify the robustness of our benchmark.
Textual Approximation of Image and Audio.
For most of the existing MLLMs that only support two input modalities (image-text or audio-text), we build up a simulated evaluation setting allowing us to explore the potential of these models to become omni-languge models in the future. We use the audio transcript annotated by human as the alternative of the audios to enable the evaluation on vision-language models. Regarding the audio-language models, we generate high-quality detailed captions of images automatically with a state-of-the-art VLM, InternVL-2-76B. In such an approximated evaluation setting, models go through the same process of inference and metric calculation as the vanilla one with textual alternatives of images or audios.
Table 2: Overall Omini-Undesratnding Results on Baseline Omni-Language Models. The overall, image-ablated and audio-ablated results on all samples are put at the first three columns. The accuracy calculated by different audio types the last three columns.
Accuracy $\uparrow$ All Audio Types Speech Sound Event Music
Input Context Image & Audio Audio Image Image & Audio
AnyGPT (7B) 2.71% 2.36% 1.23% 3.11% 1.89% 1.89%
video-SALMONN (13B) 11.30% 11.56% 11.38% 13.49% 9.06% 0.94%
UnifiedIO2-large (1.1B) 22.68% 24.69% 24.52% 21.92% 23.02% 27.36%
UnifiedIO2-xlarge (3.2B) 20.40% 24.78% 24.34% 22.96% 15.09% 15.09%
UnifiedIO2-xxlarge (6.8B) 23.29% 27.06% 25.04% 25.29% 19.62% 17.92%
Gemini-1.5-Pro 47.56% 38.53% 34.68% 47.90% 46.36% 48.11%
Reka-core-20240501 36.10% 35.07% 34.39% 38.06% 29.81% 37.74%
5 Findings
5.1 Results on Existing Omni-Language Models
Table 2 demonstrates that open-source omni-language model (OLM) baselines barely surpass random guessing accuracy across various settings. The lack of significant performance improvement in models with larger text backbones suggests that current open OLMs struggle to effectively interpret multimodal contexts and comprehend relationships between elements.
Despite overall poor performance, open-source baselines generally exhibit higher accuracy on speech audio, indicating a potential bias towards speech data. In contrast, Gemini-1.5-Pro and Reka-core-20240501, the two available proprietary models evaluated in this tri-modal setting, shows more promising results. Regarding the scores across audio types, the Gemini-1.5-Pro shows a more balanced performances while Reka-core-20240501 showing a lag on modeling the sound events. Moreover, the comparison of Gemini-1.5-Pro’s performance across full input context and ablated settings (image-text and audio-text) suggests that it effectively leverages information from all modalities to enhance its reasoning capabilities. While it demonstrates superior overall performance and balanced accuracy across audio types compared to open-source alternatives, its accuracy remains below 50%.
These findings underscore the challenging nature of OmniBench and highlight substantial room for improvement in multi-modal reasoning tasks. We anticipate the development of more competitive models on our benchmark in the near future, which will further advance the field of multi-modal AI.
Table 3: Results on Textual Audio Approximation Experiments. All the audios are represented in text transcript. The results are divided into groups of vision-language models and omni-models. We use the text transcript to approximate the audios in this setting.
Input Context Image & Audio Transcript Audio Transcript Image
InternVL-2-2B 30.04% 19.61% 23.29%
InternVL-2-8B 35.38% 24.69% 27.76%
InternVL-2-26B 34.68% 22.77% 26.80%
InternVL-2-40B 35.03% 24.34% 26.53%
Qwen2-VL-Chat-2B 30.21% 22.33% 30.74%
Qwen2-VL-Chat-7B 33.45% 24.96% 30.30%
Deepseek-VL-Chat-7B 27.67% 20.75% 23.12%
Idefics2-8B 7.09% 11.65% 1.66%
Mantis-Idefics-8B 4.29% 2.36% 0.26%
LLaVA-OneVision-0.5B 33.01% 30.56% 30.65%
LLaVA-OneVision-7B 33.27% 25.48% 23.29%
Cambrian-8B 30.65% 23.56% 25.57%
Cambrian-13B 35.64% 25.83% 31.26%
Cambrian-34B 28.46% 20.14% 22.24%
XComposer2-4KHD (7B) 30.91% 20.75% 24.87%
GPT4-o (0513) 57.62% 46.58% 43.73%
GPT4-o (0806) 53.46% 48.25% 38.35%
GPT4-o-mini 48.90% 39.58% 35.33%
Gemini-1.5-Pro 49.91% 31.79% 34.68%
Reka-core-20240501 48.73% 36.33% 34.39%
Claude-3.5-Sonnet 59.46% 37.48% 45.18%
GPT-4V-Preview 48.19% 44.75% 37.33%
GPT-4V-0409 42.49% 46.06% 37.72%
UnifiedIO2-large (1.1B) 30.56% 31.00% 24.52%
UnifiedIO2-xlarge (3.2B) 27.32% 32.57% 24.34%
UnifiedIO2-xxlarge (6.8B) 33.63% 27.85% 25.04%
Table 4: Results on Textual Image Approximation Experiments. All the images are represented in text caption. The results are divided into groups of audio-language models and omni-models.
Accuracy $\uparrow$ All Audio Types Speech Sound Event Music
Input Context Image Caption & Audio Audio Image Caption Image Caption & Audio
LTU (7B) 23.38% 26.88% 23.29% 25.68% 19.25% 16.98%
LTU-AS (7B) 28.90% 25.31% 27.06% 29.31% 29.81% 23.58%
Mu-LLaMA (7B) 23.73% 25.74% 24.52% 22.70% 26.04% 25.47%
MusiLingo-long-v1 29.77% 25.22% 27.41% 28.79% 30.57% 34.91%
audio-SALMONN (13B) 26.18% 23.12% 30.39% 26.85% 24.91% 24.53%
Qwen-Audio-Chat 29.95% 28.72% 28.28% 27.89% 32.83% 37.74%
Qwen2-Audio-7B-Instruct 43.35% 37.72% 39.67% 43.19% 44.91% 40.57%
Audio-Flamingo (1.3B) 26.80% 25.74% 24.87% 25.29% 30.94% 27.36%
Gemini-1.5-Pro 43.93% 38.53% 30.21% 46.17% 39.62% 52.83%
Reka-core-20240501 34.57% 35.07% 32.94% 35.09% 33.21% 34.29%
UnifiedIO2-large (1.1B) 27.85% 24.69% 28.81% 27.50% 30.19% 24.53%
UnifiedIO2-xlarge (3.2B) 26.80% 24.78% 29.07% 28.53% 24.53% 19.81%
UnifiedIO2-xxlarge (6.8B) 28.46% 27.06% 26.44% 27.89% 33.96% 18.87%
5.2 Textual Approximation on Images and Audios
As the absence of effective OLMs on the OmniBench, we further introduce the text alternatives of the images ($I^{\prime}$) and audios ($A^{\prime}$) to embrace more dual-modal image-text or audio-text language models to analyze the research progress of the topic. Using the textual approximations of the multimodal contexts, we evaluate the baselines and put the results using audio transcripts, image captions, and both of them in Table 3, Table 4 and Table 5 correspondingly.
Performance Changes of Open OLMs.
We select the UnifiedIO-2 series OLMs to conduct the textual approximation experiments due to their relatively robust performances in the vanilla OmniBench evaluation setting suggested in Table 2. Compared with the vanilla setting, UnifiedIO-2 models shows different levels of performance gains in this settings, i.e. averagely $8.38\%$, $5.58\%$ and $9.4\%$ when replacing the audio, image, and both of them with corresponding texts. This indicates the shortcoming of existing OLMs on modeling the image and audio (especially the latter), and a defect of interpreting these multimodal information correctly in the instruction context.
Performances of Dual-modal MLLMs.
In the setting of using text as the alternatives of audios and images, the VLMs show ALMs a generally better results (Table 3 vs Table 4). This could be caused by : 1) more available research resources including datasets and methods are allocated to image-text models; 2) the general audio are naturally harder to annotate. If the textual alternatives of images and audios could regarded as in equal quality, it seems to be easier for the future researcher to train an omni-language model from exisiting VLMs rather than ALMs.
Table 5: Results on Pure Textual Approximation for Both Image and Audio. All the images and audios are represented in texts. The results at the second and third column are taken from the corresponding models in Table 3 and Table 4.
Input Context Image Caption & Audio Transcript Audio Transcript Image Caption
LTU (7B) 22.77% 23.91% 22.59%
LTU-AS (7B) 29.16% 26.01% 26.09%
Mu-LLaMA (7B) 26.27% 25.74% 24.52%
MusiLIngo-long-v1 (7B) 28.46% 30.91% 27.41%
audio-SALMONN-13B 28.46% 27.93% 30.39%
Qwen-Audio-Chat 31.87% 34.15% 28.28%
Qwen2-Audio-7B-Instruct 47.02% 39.05% 39.67%
Audio-Flamingo (1.3B) 23.99% 25.92% 24.87%
InternVL-2-8B 32.84% 24.69% 27.67%
InternVL-2-40B 30.30% 24.34% 25.48%
Cambrian-13B 32.57% 25.83% 27.15%
GPT4-o (0513) 60.60% 46.58% 38.79%
GPT4-o (0806) 56.76% 48.25% 36.69%
GPT4-o-mini 51.41% 39.58% 33.86%
Gemini-1.5-Pro 45.27% 31.79% 30.21%
Claude-3.5-Sonnet 57.44% 37.48% 41.33%
Reka-core-20240501 42.23% 36.33% 32.94%
GPT-4V-Preview 45.51% 44.75% 33.61%
GPT-4V-0409 42.05% 46.06% 31.63%
UnifiedIO2-large (1.1B) 29.77% 31.00% 28.81%
UnifiedIO2-xlarge (3.2B) 31.70% 32.57% 29.07%
UnifiedIO2-xxlarge (6.8B) 33.10% 27.85% 26.44%
Pure Textual Evaluation.
The results in Table 5 reveal interesting patterns in model performance when both image and audio inputs are replaced with textual descriptions. Notably, the majority of models demonstrate improved accuracy when processing textual representations of multimodal data compared to their performance on either image captions or audio transcripts alone. This suggests that these models are better equipped to integrate information from multiple textual sources rather than handling raw multimodal inputs. For instance, Qwen2-Audio-7B-Instruct shows a significant jump in accuracy from 39.05% (audio transcript only) and 39.67% (image caption only) to 47.02% when given both textual representations. Similarly, proprietary models like GPT4-o and Claude-3.5-Sonnet exhibit substantial gains, with GPT4-o (0513) achieving an impressive 60.60% accuracy in the pure textual setting.
6 Conclusion and Future Study
The proposed novel multimodal benchmark, OmniBench, reveals that current open-source multimodal large language models struggle with simultaneous processing of visual, acoustic, and textual inputs. We observed a general bias towards speech audio and superior performance of vision-language models over audio-language models when using textual approximations. These findings underscore the need for more appropriate architecture designs for multimodal integration, diverse datasets for training, and techniques to reduce modality bias. OmniBench serves as a crucial tool for guiding advancements in multimodal language models, driving progress towards more advanced and versatile models towards human-like multimodal understanding and reasoning.